# An Information-Theoretic Framework for Receiver Quantization in Communication

### v1

@@ Abstract
We investigate information-theoretic limits and design of communication under receiver quantization. Unlike most existing studies that focus on low-resolution quantization, this work is more focused on the impact of weak nonlinear distortion due to resolution reduction from high to low. We consider a standard transceiver architecture, which includes independent and identically distributed (i.i.d.) complex Gaussian codebook at the transmitter, and a symmetric quantizer cascaded with a nearest neighbor decoder at the receiver. Employing the generalized mutual information (GMI), an achievable rate under general quantization rules is obtained in an analytical form, which shows that the rate loss due to quantization is $\log\left(1+\gamma\mathsf{SNR}\right)$, where $\mathsf{SNR}$ is the signal-to-noise ratio at the receiver front-end, and $\gamma$ is determined by thresholds and levels of the quantizer. Based on this result, the performance under uniform receiver quantization is analyzed comprehensively. We show that the front-end gain control, which determines the loading factor (normalized one-sided quantization range) of quantization, has an increasing impact on performance as the resolution decreases. In particular, we prove that the unique loading factor that minimizes the mean square error (MSE) of the uniform quantizer also maximizes the GMI, and the corresponding irreducible rate loss is given by $\log\left(1+\mathsf{mmse}\cdot\mathsf{SNR}\right)$, where $\mathsf{mmse}$ is the minimum MSE normalized by the variance of quantizer input, and is equal to the minimum of $\gamma$. A geometrical interpretation for the optimal uniform quantization at the receiver is further established. Moreover, by asymptotic analysis, we characterize the impact of biased gain control, including how small rate losses decay to zero and achievable rate approximations under large bias. From asymptotic expressions of the optimal loading factor and $\mathsf{mmse}$, approximations and several “per-bit rules” for performance are also provided. Finally we discuss more types of receiver quantization and show that the consistency between achievable rate maximization and MSE minimization does not hold in general.
Index Terms:
Achievable rate, analog-to-digital converter, Gaussian channel, generalized mutual information, nearest neighbor decoding rule, mean square error, MMSE, transceiver design, uniform quantization.
@@ I-A Related Work on Quantization at Communication Receivers
In vector (MIMO) channels, the quantization loss in achievable rate can be very small at low-to-moderate SNR even if $1$~$3$-bit output quantization is used. This fact was shown in information theoretic studies [9, 10, 11] by numerical examples for full-rank channels with multiplexing gains $2$~$4$. Although the constrained capacity is still unknown even in the one-bit case, recent information theoretic studies have provided various results on MIMO systems with coarse output quantization [12, 13, 14, 15, 16, 17, 18, 19], which show that proper transceiver design is critical in realizing efficient communication in such systems. In particular, for one-bit output quantization, it was shown that the high-SNR capacity grows linearly with the rank of channel [13], while the low-SNR asymptotic power loss of $\pi/2$ still exists in the case of vector channel [12].
@@
In light of these positive theoretical results, performance and design of wireless systems with low-resolution quantization have been extensively studied in recent years; see, e.g., [20, 21, 22, 23, 24, 25, 26]. The most common approach therein, however, is not information-theoretic (since exact evaluation of mutual information can be difficult). Instead, achievable rate estimation based on the additive quantization noise model (AQNM) has been widely used, which comes from Bussgang-like decomposition [27, 28, 29]. Results in [20, 12, 13, 14] suggested that such estimation approximates the mutual information well at low SNR, but becomes inaccurate at high SNR.
@@
Although mutual information is a fundamental performance measure, for communications under transceiver nonlinearity it has limited operational meaning, in the sense that the decoder that achieves the predicted rate can be too complex to implement, while that rate is not necessarily achievable by a standard transceiver architecture designed without considering nonlinearity because the decoder is typically mismatched to the nonlinear channel. In [30, 31], a more meaningful performance measure that takes decoding rule into account, namely the generalized mutual information (GMI) [32], has been adopted, yielding analytical expressions of the achievable rate under output quantization and nearest neighbor decoding rule. Under a given (possibly mismatched) decoder, the GMI determines the highest rate below which the average probability of error, averaged over a given i.i.d. codebook ensemble, converges to zero as the block length $N$ grows without bound, and it is thus a lower bound on mismatch capacity [32, 33]. The GMI has been applied in various scenarios for performance evaluation, including the bit-interleaved coded modulation (BICM) [34], fading channels [35, 36, 37], and nonlinear fiber-optic channels [38]. In fact, the rate estimation based on the AQNM (or Bussgang-like decomposition) is consistent with the GMI for scalar channel under Gaussian input and nearest neighbor decoding [30]; see [39] for more discussions. However, GMI analyses also show that the AQNM-based estimation is not accurate in general. For example, it may overestimate the achievable rate in multiantenna systems [40].
@@ I-B Related Work in Quantization Theory
The rich theory of quantization was surveyed comprehensively in [41], in which two well established asymptotic theories were emphasized. The first is Shannon’s information theoretic approach (rate distortion theory [42]), which places quantization in the framework of lossy source coding and focuses on the high-dimension regime, thereby shedding light on vector quantization. The second is the asymptotic quantization theory, which sheds light on quantizer design in the high-resolution regime. The asymptotic quantization theory is more relevant to receiver quantization in communication which typically does not employ coding or vector quantization. Although not applicable directly to receiver quantization, some classical results in asymptotic quantization theory, especially those for uniform scalar quantization, are reviewed here for comparison.
A basic result known in [43] and [44] (rigorously proved in [45]) states that, for a high-resolution uniform quantizer with step size $\ell$, the mean square error (MSE) can be approximated by $\ell^{2}/12$. This yields the “$6$-dB-per-bit rule” that each additional bit in resolution reduces the MSE by $6.02$ dB. The rule reflects the impact of step size that causes granular distortion; but it ignores overload distortion due to finite quantization range. The interplay between these two types of distortion determines how the optimal quantization range scales with the resolution. The scaling law has been characterized in the seminal work [46] for several types of input densities. Take the Gaussian source as example. In [46] it has been shown that, for the optimal $2K$-level uniform quantization that minimizes the MSE, 1) the loading factor [^2] The loading factor of a quantizer is the one-sided width of its quantization range normalized by the standard deviation of the input [47]. scales like $2\sqrt{\ln(2K)}$ (cf. the conventional “four-sigma” rule of thumb [44, 48, 47]), and 2) the granular distortion dominates and the overload distortion is asymptotically negligible. Further properties of the uniform quantization have been analyzed in [49, 50, 51]. For quantization at communication receivers, we need parallel results to characterize the optimal loading factor, which is essential for the design of AGC. We note that, although the importance of AGC design in the presence of output quantization has been recognized for a long time [6, 29], it was only investigated by numerical results in several works; see, e.g., [9, 10, 52, 53, 22].
@@ I-C1 Problem
In this paper, we consider communications in the presence of output quantization, and focus on the effect of resolution reduction on the performance of a standard transceiver architecture designed without considering that effect. Rather than considering only low-resolution output quantization ($1$~$3$ bits) as in most existing studies, we consider the entire region of resolution, especially the transition from high resolution (typically $8$~$12$ bits or more so that performance loss is negligible) to low resolution. Since in the transition only weak to moderate nonlinearity is introduced, it is natural to keep the transceiver architecture unchanged rather than rebuild it. The considered standard transceiver architecture includes independent and identically distributed (i.i.d.) complex Gaussian codebook at the transmitter, and a uniform output quantizer cascaded with a (weighted) nearest neighbor decoder at the receiver, where the loading factor of the quantizer can be adjusted by gain control. See Sec. II for details. We choose such an architecture in view of the following facts.
  • When the impact of quantization is negligible, this architecture is capacity-achieving in several important channel models, such as the AWGN channel and the flat-fading channel with Gaussian noise and channel state information at the receiver [55]. It is also a very robust architecture in general noisy channels [56]. This architecture has been adopted in various performance evaluation problems, e.g., [56, 35, 36] for linear channels with fading and [30, 38, 57] for nonlinear channels.
@@ I-C3 Summary of Contribution
  • The uniform receiver quantization is further studied by asymptotic analysis. In Sec. V, the impact of biased gain control is characterized by asymptotic behaviors and approximations of $I_{\textrm{GMI}}(L)$. Approximations of $I_{\textrm{GMI}}(L)$ for overload region ($L<L^{*}$) and underload region ($L>L^{*}$) of the loading factor are proposed, respectively. Specifically, in the high-resolution regime, we characterize the loading loss in achievable rate, showing that i) the loading loss due to overload distortion decays exponentially as the loading factor increases, and ii) the loading loss due to granular distortion decays quadratically as the step size decreases. In Sec. VI, focusing on the optimal uniform quantization, we provide a new approximation of $L^{*}$, an approximation of $I_{\textrm{GMI}}(L^{*})$, and several per-bit rules for performance metrics such as saturation rates and irreducible rate loss;
  • In Sec. VII, for general quantization rules, we illustrate that the consistency between achievable rate maximization and MSE minimization does not necessarily hold. We also discuss the possible gain of introducing more types of quantization for communication receiver.
@@ II-A A Standard Transceiver Architecture
The transceiver architecture we consider is shown in Fig. 1. For a code rate $R$ bits/channel use (c.u.), a message is selected uniformly randomly from the index set =$\mathcal{M}=\{1,2,...,\lceil 2^{NR}\rceil\}$. If a message $m$ is selected, then the encoder maps it to a codeword $[X_{1}(m),...,X_{N}(m)]$ of block length $N$, which is generated according to a product complex Gaussian distribution $\mathcal{N}_{\mathbb{C}}(0,\sigma_{x}^{2}\mathbf{I}_{N})$. During transmission, each transmitted symbol is scaled by a channel gain $h\in\mathbb{C}$ which remains constant over the transmission duration of a codeword. The scaled symbols are corrupted by i.i.d. complex Gaussian noise at the receiver front-end before quantization. Then the channel output after quantization is given by
@@
where $n=1,...,N$, $q(\cdot)$ denotes the quantizer which introduces nonlinear distortion, $g\in\mathbb{R}^{+}$ is a gain-control factor, and
@@ II-A1 Nearest Neighbor Decoding Rule
That is, it selects a message corresponding to the codeword (scaled by a parameter $a$) with the minimum Euclidean distance to the received vector $[Y_{1},...,Y_{N}]$. [^3] Some more general forms of nearest neighbor decoding rule can be found in, e.g., [56, 35, 36]. A more recent work including a detailed review and some generalizations is [39].
@@ II-A2 Symmetric Quantizer and Uniform Quantizer
Let the quantizer in (4) be symmetric with $2K$ representation points (levels) {±$\{\pm y_{1},...,\pm y_{K}\}$ and normalized thresholds $\{0,\pm l_{1},...,\pm l_{K-1}\}$, where $y_{k}\geq 0$ and $l_{k}\geq 0$ for $k=1,...,K$. Then its resolution (bit-width) is $b=\log_{2}{2K}$ bits, which typically satisfies $b\in\mathbb{Z}^{+}$. Let $V\in\mathbb{R}$ be the input to be quantized with standard deviation $\sigma_{v}$. For both quantizers in (4) we have $\sigma_{v}=\sqrt{(|h|^{2}\sigma_{x}^{2}+\sigma^{2})/2}$. Then for the input $V$, the output of the quantizer is
@@
where $\ell$ is the step size. Thus, we define its quantization range or support as $[-K\ell,K\ell]$. Then the loading factor or support limit of the uniform quantizer is $L=K\ell$.
@@ II-B An Achievable Rate Formula from Generalized Mutual Information
Proposition 1 [30]: For a memoryless SISO channel $X\to Y$ with transition probability $p_{Y|X}(y|x)$ and nearest neighbor decoding rule (7), where $X,Y\in\mathbb{C}$ and $\mathrm{Var}(X)=\sigma_{x}^{2}$, the maximum GMI under i.i.d. complex Gaussian codebook is given by
@@
Let $\sigma_{y}$ be the standard deviation of $Y$. We note that when $\mathrm{E}[Y]=0$ (e.g., when the quantizer is symmetric, the channel output in (4) has zero mean) we have $\Delta=|\rho_{XY}|^{2}$, where
@@
is the Pearson correlation coefficient between $X$ and $Y$. Also note that the equality in (17) holds if and only if $X\perp D$ and $D\sim\mathcal{N}_{\mathbb{C}}(0,\sigma^{2})$ (e.g., when there is no quantization in (4)), so that the channel $X\to Y$ reduces to the AWGN channel, the nearest neighbor decoding rule is optimal, and the GMI equals to the mutual information (also the channel capacity) $\log\left(1+\frac{\sigma_{x}^{2}}{\sigma^{2}}\right)$.

### v2

@@ Abstract
We investigate information-theoretic limits and design of communication under receiver quantization. Unlike most existing studies that focus on low-resolution quantization, this work is more focused on the impact of weak nonlinear distortion due to resolution reduction from high to low. We consider a standard transceiver architecture, which includes independent and identically distributed (i.i.d.) complex Gaussian codebook at the transmitter, and a symmetric quantizer cascaded with a nearest neighbor decoder at the receiver. Employing the generalized mutual information (GMI), an achievable rate under general quantization rules is obtained in an analytical form, which shows that the rate loss due to quantization is $\log\left(1+\gamma\mathsf{SNR}\right)$, where $\mathsf{SNR}$ is the signal-to-noise ratio at the receiver front-end, and $\gamma$ is determined by thresholds and levels of the quantizer. Based on this result, the performance under uniform receiver quantization is analyzed comprehensively. We show that the front-end gain control, which determines the loading factor (normalized one-sided quantization range) of quantization, has an increasing impact on performance as the resolution decreases. In particular, we prove that the unique loading factor that minimizes the mean square error (MSE) of the uniform quantizer also maximizes the GMI, and the corresponding irreducible rate loss is given by $\log\left(1+\mathsf{mmse}\cdot\mathsf{SNR}\right)$, where $\mathsf{mmse}$ is the minimum MSE normalized by the variance of quantizer input, and it is equal to the minimum of $\gamma$. A geometrical interpretation for the optimal uniform quantization at the receiver is further established. Moreover, by asymptotic analysis, we characterize the impact of biased gain control, showing how small rate losses decay to zero and providing approximations for the achievable rate under large bias. From asymptotic expressions of the optimal loading factor and $\mathsf{mmse}$, approximations and several “per-bit rules” for performance are also provided. Finally we discuss more types of receiver quantization and show that the consistency between achievable rate maximization and MSE minimization does not hold in general.
@@ I-A Related Work on Quantization at Communication Receivers
In vector (MIMO) channels, the quantization loss in achievable rate can be very small at low-to-moderate SNR even if 11~33-bit output quantization is used. This fact was shown in information theoretic studies [9, 10, 11] by numerical examples for full-rank channels with multiplexing gains 22~44. Although the constrained capacity is still unknown even in the one-bit case, recent information theoretic studies have provided various results on MIMO systems with coarse output quantization [12, 13, 14, 15, 16, 17, 18, 19], which show that proper transceiver design is critical in realizing efficient communication in such systems. In particular, for one-bit output quantization, it was shown that the high-SNR capacity grows linearly with the rank of channel [13], while the low-SNR asymptotic power loss of $\pi/2$ still exists in the case of vector channel [12].
@@
In light of these positive theoretical results, performance and design of wireless systems with low-resolution quantization have been extensively studied in recent years; see, e.g., [20, 21, 22, 23, 24, 25, 26, 27]. The most common approach therein, however, is not information-theoretic (since exact evaluation of mutual information can be difficult). Instead, achievable rate estimation based on the additive quantization noise model (AQNM) has been widely used, which comes from Bussgang-like decomposition [28, 29, 30]. Results in [20, 13, 14] suggested that such estimation approximates the mutual information well at low SNR, but becomes inaccurate at high SNR. [^2] See the remark at the end of Sec. II for more discussions.
@@
Although mutual information is a fundamental performance measure, for communications under transceiver nonlinearity it has limited operational meaning, in the sense that the decoder that achieves the predicted rate can be too complex to implement, while that rate is not necessarily achievable by a standard transceiver architecture designed without considering nonlinearity because the decoder is typically mismatched to the nonlinear channel. In [31, 32], a more meaningful performance measure that takes decoding rule into account, namely the generalized mutual information (GMI) [33], has been adopted, yielding analytical expressions of the achievable rate under output quantization and nearest neighbor decoding rule. Under a given (possibly mismatched) decoder, the GMI determines the highest rate below which the average probability of error, averaged over a given i.i.d. codebook ensemble, converges to zero as the block length $N$ grows without bound, and it is thus a lower bound on mismatch capacity [33, 34]. The GMI has been applied in various scenarios for performance evaluation, including the bit-interleaved coded modulation (BICM) [35], fading channels [36, 37, 38], and nonlinear fiber-optic channels [39]. In fact, the rate estimation based on the AQNM (or Bussgang decomposition) is consistent with the GMI for scalar channel under Gaussian input and nearest neighbor decoding [31]; see [40] for more discussions.
@@ I-B Related Work in Quantization Theory
The rich theory of quantization was surveyed comprehensively in [42], in which two well established asymptotic theories were emphasized. The first is Shannon’s information theoretic approach (rate distortion theory [43]), which places quantization in the framework of lossy source coding and focuses on the high-dimension regime, thereby shedding light on vector quantization. The second is the asymptotic quantization theory, which sheds light on quantizer design in the high-resolution regime. The asymptotic quantization theory is more relevant to receiver quantization in communication which typically does not employ coding or vector quantization. Some classical results in asymptotic quantization theory, especially those for uniform scalar quantization, are reviewed here. Although not applicable directly to receiver quantization, they can be helpful in deriving new results (e.g., Corollary 10 in Sec. IV), closely related to our work (e.g., we will show the role of mean square error (MSE) in achievable rate analysis), and comparable to our findings (e.g., a remark in Sec. VII-A).
A basic result known in [44] and [45] (rigorously proved in [46]) states that, for a high-resolution uniform quantizer with step size $\ell$, the MSE can be approximated by $\ell^{2}/12$. This yields the “$6$-dB-per-bit rule” that each additional bit in resolution reduces the MSE by $6.02$ dB. The rule reflects the impact of step size that causes granular distortion; but it ignores overload distortion due to finite quantization range. The quantization range is often represented by a parameter called loading factor [47], which is defined as the one-sided width of quantization range normalized by the standard deviation of the input of the quantizer. The interplay between granular distortion and overload distortion determines how the optimal loading factor scales with the resolution. The scaling law has been characterized in the seminal work [48] for several types of input densities. Take the Gaussian source as example. In [48] it has been shown that, for the optimal $2K$-level uniform quantization that minimizes the MSE, 1) the loading factor scales like $2\sqrt{\ln(2K)}$ (cf. the conventional “four-sigma” rule of thumb [45, 49, 47]), and 2) the granular distortion dominates and the overload distortion is asymptotically negligible. Further properties of the uniform quantization have been analyzed in [50, 51, 52]. For quantization at communication receivers, we need parallel results to characterize the optimal loading factor, which is essential for the design of AGC. We note that, although the importance of AGC design in the presence of output quantization has been recognized for a long time [6, 30], it was only investigated by numerical results in several works; see, e.g., [9, 10, 53, 54, 22].
@@ I-C1 Problem
In this paper, we consider communications in the presence of output quantization, and focus on the effect of resolution reduction on the performance of a standard transceiver architecture designed without considering that effect. Rather than considering only low-resolution output quantization (11~33 bits) as in most existing studies, we consider the entire region of resolution, especially the transition from high resolution (typically 88~$12$ bits or more so that performance loss is negligible) to low resolution. Since in the transition only weak to moderate nonlinearity is introduced, it is natural to keep the transceiver architecture unchanged rather than rebuild it. The considered standard transceiver architecture includes independent and identically distributed (i.i.d.) complex Gaussian codebook at the transmitter, and a uniform output quantizer cascaded with a (weighted) nearest neighbor decoder at the receiver, where the loading factor of the quantizer can be adjusted by gain control. See Sec. II for details. We choose such an architecture in view of the following facts.
  • When the impact of quantization is negligible, this architecture is capacity-achieving in several important channel models, such as the AWGN channel and the flat-fading channel with Gaussian noise and channel state information at the receiver (CSIR) [56]. It is also a very robust architecture in general noisy channels [57]. This architecture has been adopted in various performance evaluation problems, e.g., [57, 36, 37] for linear channels with fading and [31, 39, 58] for nonlinear channels.
@@ I-C3 Summary of Contribution
  • The uniform receiver quantization is further studied by asymptotic analysis. In Sec. V, the impact of biased gain control is characterized by asymptotic behaviors and approximations of $I_{\textrm{GMI}}(L)$. Approximations of $I_{\textrm{GMI}}(L)$ for overload region ($L<L^{*}$) and underload region ($L>L^{*}$) of the loading factor are proposed, respectively. Specifically, in the high-resolution regime, we characterize the loading loss in achievable rate, showing that i) the loading loss due to overload distortion decays exponentially as the loading factor increases, and ii) the loading loss due to granular distortion decays quadratically as the step size decreases. In Sec. VI, focusing on the optimal uniform quantization, we provide a new approximation of $L^{*}$, an approximation of $I_{\textrm{GMI}}(L^{*})$, and several per-bit rules for performance metrics such as saturation rate and irreducible rate loss;
  • In Sec. VII, for general quantization rules, we illustrate that the consistency between achievable rate maximization and MSE minimization does not necessarily hold. In fact, except for the uniform quantizer, in our setting only two specific quantizers that possess the consistency have been found; see Sec. VII-A. We also discuss the possible gain of introducing more types of quantization for communication receiver.
@@ II-A A Standard Transceiver Architecture
The transceiver architecture we consider is shown in Fig. 1. For a code rate $R$ bits/channel use (c.u.), a message is selected uniformly randomly from the index set $\mathcal{M}=\{1,2,...,\lceil 2^{NR}\rceil\}$. If a message $m$ is selected, then the encoder maps it to a codeword $X_{1}(m),...,X_{N}(m)]$ of block length $N$, which is generated according to a product complex Gaussian distribution $\mathcal{N}_{\mathbb{C}}(0,\sigma_{x}^{2}\mathbf{I}_{N})$. During transmission, each transmitted symbol is scaled by a channel gain $h\in\mathbb{C}$ which remains constant over the transmission duration of a codeword. The scaled symbols are corrupted by i.i.d. complex Gaussian noise at the receiver front-end before quantization. Then the channel output after quantization is given by
@@
where n=1,,Nn=1,...,N, $q(\cdot)$ denotes the quantizer which introduces nonlinear distortion, $g\in\mathbb{R}^{+}$ is a gain-control factor, and
@@ II-A1 Nearest Neighbor Decoding Rule
That is, it selects a message corresponding to the codeword (scaled by a parameter $a$) with the minimum Euclidean distance to the received vector $Y_{1},...,Y_{N}]$. [^3] Some more general forms of nearest neighbor decoding rule can be found in, e.g., [57, 36, 37]. A more recent work including a detailed review and some generalizations is [40].
@@ II-A2 Symmetric Quantizer and Uniform Quantizer
Let the quantizer in (4) be symmetric with $2K$ representation points (levels) $\{\pm y_{1},...,\pm y_{K}\}$ and normalized thresholds $\{0,\pm l_{1},...,\pm l_{K-1}\}$, where $y_{k}\geq 0$ and $l_{k}>0$ for k=1,,Kk=1,...,K, and there are at least two nonzero levels. Then its resolution (bit-width) is $b=\log_{2}{2K}$ bits, which typically satisfies $b\in\mathbb{Z}^{+}$. Let $V\in\mathbb{R}$ be the input to be quantized with standard deviation $\sigma_{v}$. The inputs of both quantizers in (4) have the same standard deviation, namely $\sigma_{v}=\sqrt{(|h|^{2}\sigma_{x}^{2}+\sigma^{2})/2}$. Then for the input $V$, the output of the quantizer is
@@
where $\ell$ is the step size. Thus, we define its quantization range or support as $K\ell,K\ell]$. Then the loading factor or support limit of the uniform quantizer is $L=K\ell$.
@@ II-B An Achievable Rate Formula from Generalized Mutual Information
Proposition 1 [31]: For a memoryless scalar channel $X\to Y$ with transition probability $p_{Y|X}(y|x)$ and nearest neighbor decoding rule (7), where $X,Y\in\mathbb{C}$ and $\mathrm{Var}(X)=\sigma_{x}^{2}$, the maximum GMI under i.i.d. complex Gaussian codebook is given by
@@
We note that, in Proposition 1, if $\mathrm{E}[Y]\neq 0$, then a pre-processing of the channel output as $Y\to\tilde{Y}=Y-\mathrm{E}[Y]$ improves the GMI (14). When $\mathrm{E}[Y]=0$ has been satisfied (e.g., the output $Y_{n}$ in (4) when $q(\cdot)$ is a symmetric quantizer), we have $\Delta=|\rho_{XY}|^{2}$, where
@@
is the Pearson correlation coefficient between $X$ and $Y$, where $\sigma_{y}$ is the standard deviation of $Y$. Also note that the equality in (17) holds if and only if $X\perp D$ and $D\sim\mathcal{N}_{\mathbb{C}}(0,\sigma^{2})$ (e.g., when there is no quantization in (4)), so that the channel $X\to Y$ reduces to the AWGN channel, the nearest neighbor decoding rule is optimal, and the GMI equals to the mutual information (also the channel capacity) $\log\big(1+\frac{\sigma_{x}^{2}}{\sigma^{2}}\big)$.
Remark (Bussgang decomposition and AQNM): Proposition 1 and its proof provide an information-theoretic interpretation to the achievable rate estimation based on Bussgang decomposition $Y=\alpha X+D$. The estimated rate is typically given in terms of an effective SNR (or signal to noise-and-distortion ratio) as
$R=\log\left(1+\mathsf{SNR}_{\mathrm{e}}\right),$ (19)
where
$\mathsf{SNR}_{\mathrm{e}}=\sup\limits_{a\in\mathbb{R}}\frac{\mathrm{E}[|aX|^{2}]}{\mathrm{E}[|Y-aX|^{2}]}=\frac{\mathrm{E}[|\alpha X|^{2}]}{\mathrm{E}[|D|^{2}]},$ (20)
and $\alpha$ is called the Bussgang gain. It is straightforward to verify that $\mathsf{SNR}_{\mathrm{e}}=\Delta/(1-\Delta)$, so that the rate (19) is equal to the GMI expression (14) in Proposition 1. In [40] this interpretation is extended to a more general setting, which includes memoryless channel state, imperfect CSIR, and vector (rather than scalar) channel output (but still assume i.i.d. Gaussian input). In brief, when the receiver linearly combines elements of channel output into a scalar for nearest neighbor decoding, the linear combiner that maximizes the GMI is exactly the linear minimum MSE (LMMSE) estimator of the channel input upon observing the channel output, and the maximum GMI is consistent with the achievable rate estimation based on Bussgang decomposition; see [References, Sec. IV-C] for details. [^5] In the description in [References, Sec. IV-C], it was incorrectly stated that the distortion is zero-mean. This may not hold for general nonlinear channels. But this does not affect the conclusion since in the effective SNR (cf. (20)), the distortion power is defined as the second-order moment, instead of the variance, of distortion. (Discussion with Yuhao Liu of Tsinghua University is acknowledged) Since the interpretation relies on i.i.d. Gaussian codebook, nearest neighbor decoding rule, and certain receiver processing, the following issues should be noted.
  • To achieve accurate rate estimation, the AQNM should be derived rigorously according to the Bussgang decomposition. In fact, this is not always satisfied in the literature. For example, in multiantenna systems, treating the residual noise in the AQNM as i.i.d. Gaussian vector (omitting the correlation among its elements) leads to inaccurate results, as pointed out in, e.g., [20], [41] (which includes an example of significant overestimate of achievable rate), and [29].
  • The GMI does not necessarily approach the mutual information with the same input distribution in general channels. Thus, the AQNM should not be expected to provide an accurate estimate of the mutual information, especially at high SNR, as observed in [20, 13, 14].
  • No information-theoretic foundation has been established for the generalized Bussgang decomposition [29] for non-Gaussian input signals.